OpenStudy (anonymous):

Identify the GCF of 12a^4b^3 + 8a^3b^2?

OpenStudy (saifoo.khan):

Try finding the commons.

OpenStudy (anonymous):

two terms?

OpenStudy (anonymous):

First seperate them 12a^4b^3 =(2)(2)(3)(a)(a)(a)(a)(b)(b)(b) 8a^3b^2 =(2)(2)(2)(a)(a)(a)(b)(b) Take terms that at least one of them have GCF=(2)(2)(2)(3)(a)(a)(a)(a)(b)(b)(b)